4. Problems and Applications Q5

Economists in Funlandia, which has a closed economy, have collected the following information about the economy for a particular year:

Y = 12,500
C = 9,000
T = 2,100
G = 2,200

The economists also estimate that the investment function is:

I = 2,000100r

where r is the country’s real interest rate, expressed as a percentage.

Complete the following table by calculating private saving, public saving, national saving, investment, and the equilibrium real interest rate.

ComponentAmount
Private Saving

1,400

Public Saving

-100

National Saving

1,300

Investment

1,300

Equilibrium Real Interest Rate

7%
Answer #1

Private saving is the amount of income that households have left after paying their taxes (T) and paying for their consumption (C):

SPrivate = YTC

 = 12,5002,1009,000

 = 1,400

Public saving is the amount of tax revenue that the government has left after paying for its spending (G):

SPublic = TG

 = 2,1002,200

 = −100

National saving is the sum of private and public saving:

National Saving = SPrivate+SPublic

 = 1,400+−100

 = 1,300

Because investment must equal national saving in a closed economy, investment is also $1,300.

You can find the equilibrium interest rate by setting investment equal to 1,300 and solving for r:

I = 2,000100r
1,300 = 2,000100r
100r = 2,0001,300
100r = 700
r = 7%


answered by: MARK NJOROGE
Answer #2

SOLUTION :


Y = 12500 . C = 9000, T = 2100, G = 2300 and I = 2000 - 100r 


Private saving 

= Total income - consumption - taxes 

= Y - C - T

= 12500 - 9000 - 2100

= 1400


Public saving 

= Govt. revenue - Govt. expenses

= T - G 

= 2100 = 2200

= - 100 


National saving 

= Private saving + Public saving

= 1400 + (-100)

= 1300 


Investment = 2000 - 100r

=>  National saving = 2000 - 100r

= > 1300 = 2000 - 100r

So,

 r = interest rate 

= (2000 - 1300) / 100 

= 7%


Hence, the table will be as under :


Details                          Value


Private saving               1400


Public saving                 - 100


National saving             1300


Interest rate                     7%

answered by: Tulsiram Garg